# MobiLLM: An Agentic AI Framework for Closed-Loop Threat Mitigation in 6G Open RANs

### v1

@@ I Introduction
The current research landscape reveals two parallel streams of work in O-RAN: passive security detection and non-security-focused automation. On one hand, significant research has concentrated on threat detection. This includes various xApps designed to identify network anomalies [23] and malicious activities such as the presence of rogue UEs and base stations [22, 25, 10]. A critical limitation of these solutions, however, is their passive nature; they are designed to report threats but lack the capability to automatically respond or mitigate them. On the other hand, numerous closed-loop xApps and rApps have been developed for network automation and performance optimization tasks like network slicing and resource management [20, 14, 5, 28]. These systems, while demonstrating the potential of automation, were not created to address cybersecurity challenges. This leaves a crucial gap for a framework that can unify intelligent detection with proactive, closed-loop security response.
@@
  • Development of MobiLLM, the first multi-tiered agentic framework for 5G threat analysis, planning and response.
  • Evaluation of MobiLLM with a suite of five real world 5G threat scenarios. We release associated RRC protocol traces and initial incident reports for each threat.
@@ II-A Threat Analysis Agent
The Threat Analysis Agent serves as the initial triage system. Its primary input is not just a simple prompt, but a detected threat or abnormal event, which could originate from network monitoring tools, an O-RAN xApp/rApp, or a natural language query from a human operator. Upon receiving an alert, the agent’s first moves beyond the surface-level description by actively inspecting the event and correlating it with related network data, querying sources such as performance metrics, RAN-level data, and logs. This process allows the agent to analyze the context, evaluate the potential risk, and distinguish a genuine security threat from a benign network fault or false positive. If a credible threat is identified, the agent generates a structured, machine-readable threat report containing the event summary, affected components, and an initial risk assessment. If no threat is detected, it logs the analysis and closes the event, ensuring that downstream agents do not expend unnecessary computational resources.
@@ II-D Implementation Details
    Control APIs: These functions enable the RAN configuration tuning agent to to perform an action in a human-supervised control loop. An agent can fetch the active configuration of a CU, suggest modifications, and, upon receiving human approval, apply the updated configuration and reboot the RAN to enact the changes. The scope of controllable parameters can be readily expanded by incorporating additional parameters from the O-RAN E2SM-RC (RAN Control) standard [21].
Agent Tools and Prompts. The threat analysis, threat classification, security planning and action agent interface with the knowledge bases, RAN control system, configuration tuning subroutines and system reboot functions to dynamically analyze and craft a response to an ongoing threat. The agents are free to call tools in any fashion (single-turn, multi-turn, multi-step etc.). We use ReAct style few-shot prompts that encourage the model to think, reason and act in an iterative loop. Max iterations for that loop is five.
@@ III Evaluation
Since MobiLLM represents the first known agentic architecture capable of analyzing, classifying, and autonomously responding to threats within 6G O-RAN networks, defining a robust evaluation methodology posed a significant challenge. In the absence of a standardized dataset or established evaluation protocol for this domain, we curated a focused set of five representative attacks that 6G networks inherit from prior standards (e.g., LTE and 5G NR): Blind Denial-of-Service (DoS), Base Transceiver Station (BTS) resource depletion, Downlink IMSI exposure, Uplink IMSI exposure, and Null cipher exploitation [12, 6, 11], which is reproduced end-to-end within a cellular network testbed using open-sourced software including OpenAirInterface and OpenRAN Software Community RIC.
@@ III-A Evaluation Setup
We construct a 5-case evaluation suite representative of the aforementioned 6G attacks. For each threat we record (i) A threat scenario describing the threat and expected observables; (ii) Curated telemetry (RRC and NAS traffic logs) corresponding to the attacks; and (iii) The MITRE FiGHT [19] tactics/techniques classification associated with each threat and their corresponding mitigations. We use the scenario description and the curated telemetry as context with the initial prompt to the threat analysis agent. To evaluate the MITRE FiGHT classification performance, we ask a few cellular network experts to generate ground truth results for all the targeted attacks.
@@ IV Discussion and Future Work
MobiLLM is a promising direction towards fully automated threat incidence analysis and response. We perform all our experiments with Gemini-2.5-Flash as the core deduction engine for both analysis generation and function calling, and the distinction between agents is made through what tools they have at their disposal (geared towards a specific task). While knowledge bases and tools help ground the LLM response, further fine tuning is imperative for state of the art performance for the following reasons:
@@
    API-based off the shelf models are ephemeral and costly. Training bespoke LoRA [9] adapters enables light-weight SLMs [3] achieve better performance in agentic tasks. We plan to bootstrap the current capabilities of our Agentic architecture to create a supervised fine-tuning dataset; to train small but specialized language models specifically for the task of crafting a threat response in the context of O-RAN 5G.
@@
    Currently, we supply network facts like CU configurations and parameters at inference time, which make it slow and costly. Through supervised fine tuning, this information can be baked in the model parameters. We also plan to expand the use case to more than 5 attacks that we currently support, with tools designed to address a wide range of threats.
@@ IV-B Accessible and robust O-RAN security
We plan to integrate MobiLLM into real-world 6G deployments. As initiatives like the U.S. government’s Open Centralized Unit Distributed Unit (OCUDU) project [7] accelerate the creation of open 6G platforms, MobiLLM can be embedded as a core security layer to not only enhance network resilience and trust from the outset, but also dramatically cut down network operator training cost and time. We plan to develop more advanced security guardrails to protect MobiLLM itself from adversarial attacks like backdoor attacks [17] and prompt injection attacks [15]. This involves creating stricter, verifiable rules for its API interactions to defend against threats like prompt injection and ensure its operational integrity in mission-critical environments.
@@ Acknowledgements
This research was supported by a Small Business Innovation Research (SBIR) Phase I award N6893625C0023 from the Naval Air Warfare Center (NAWC) and the NSF convergence accelerator program under award ITE-2326882.

### v2

@@ I Introduction
The current research landscape reveals two parallel streams of work in O-RAN: passive security detection and non-security-focused automation. On one hand, significant research has concentrated on threat detection. This includes various xApps designed to identify network anomalies [22] and malicious activities such as the presence of rogue UEs and base stations [21, 24, 9, 25]. A critical limitation of these solutions, however, is their passive nature; they are designed to report threats but lack the capability to automatically respond or mitigate them. On the other hand, numerous closed-loop xApps and rApps have been developed for network automation and performance optimization tasks like network slicing and resource management [19, 13, 5, 28]. These systems, while demonstrating the potential of automation, were not created to address cybersecurity challenges. This leaves a crucial gap for a framework that can unify intelligent detection with proactive, closed-loop security response.
@@
  • Development of MobiLLM, the first multi-tiered agentic framework for cellular threat analysis, planning, and response  [^1] MobiLLM has been released at https://github.com/5GSEC/MobiLLM.
  • Evaluation of MobiLLM with a suite of five real-world 5G threat scenarios. We release associated RRC protocol traces and initial incident reports for each threat.
@@ II-A Threat Analysis Agent
The Threat Analysis Agent serves as the initial triage system. Its primary input is not just a simple prompt, but a detected threat or abnormal event, which could originate from network monitoring tools, an O-RAN xApp/rApp, or a natural language query from a human operator. Upon receiving an alert, the agent’s first moves is beyond the surface-level description by actively inspecting the event and correlating it with related network data, querying sources such as performance metrics, RAN-level data, and logs. This process allows the agent to analyze the context, evaluate the potential risk, and distinguish a genuine security threat from a benign network fault or false positive. If a credible threat is identified, the agent generates a structured, machine-readable threat report containing the event summary, affected components, and an initial risk assessment. If no threat is detected, it logs the analysis and closes the event, ensuring that downstream agents do not expend unnecessary computational resources.
@@ II-D Implementation Details
    Control APIs: These functions enable the RAN configuration tuning agent to perform an action in a human-supervised control loop. An agent can fetch the active configuration of a CU, suggest modifications, and, upon receiving human approval, apply the updated configuration and reboot the RAN to enact the changes. The scope of controllable parameters can be readily expanded by incorporating additional parameters from the O-RAN E2SM-RC (RAN Control) standard [20].
Agent Tools and Prompts. The threat analysis, threat classification, security planning and action agent interface with the knowledge bases, RAN control system, configuration tuning subroutines and system reboot functions to dynamically analyze and craft a response to an ongoing threat. The agents are free to call tools in any fashion (single-turn, multi-turn, multi-step, etc.). We use React-style few-shot prompts that encourage the model to think, reason, and act in an iterative loop. Max iterations for that loop is five.
@@ III Evaluation
Since MobiLLM represents the first agentic architecture capable of analyzing, classifying, and autonomously responding to threats within 6G O-RAN networks, defining a robust evaluation methodology posed a significant challenge. In the absence of a standardized dataset or established evaluation protocol for this domain, we curated a focused set of five representative attacks that 6G networks inherit from prior standards (e.g., LTE and 5G NR): Blind Denial-of-Service (DoS), Base Transceiver Station (BTS) resource depletion, Downlink IMSI exposure, Uplink IMSI exposure, and Null cipher exploitation [11, 6, 10], which were reproduced end-to-end within a testbed using open-sourced software including OpenAirInterface and OpenRAN Software Community RIC.
@@ III-A Evaluation Setup
We construct a 5-case evaluation suite representative of the aforementioned 6G attacks. For each threat we record (i) A threat scenario describing the threat and expected observables; (ii) Curated telemetry (RRC and NAS traffic telemetry) corresponding to the attacks; and (iii) The MITRE FiGHT [18] tactics classification associated with each threat and their corresponding mitigations. We use the scenario description and the curated telemetry as context with the initial prompt to the threat analysis agent. To evaluate the threat classification performance, we ask a few cellular network experts to generate ground truth results for all the targeted attacks.
@@ IV Discussion and Future Work
MobiLLM is a promising direction towards fully automated threat incidence analysis and response. We perform all our experiments with Gemini-2.5-Flash as the core deduction engine for both analysis generation and function calling, and the distinction between agents is made through what tools they have at their disposal (geared towards a specific task). While knowledge bases and tools help ground the LLM response, further fine tuning is imperative for state-of-the-art performance for the following reasons:
@@
    API-based off-the-shelf models are ephemeral and costly. Training bespoke LoRA [8] adapters enables light-weight SLMs [3] to achieve better performance in agentic tasks. We plan to bootstrap the current capabilities of our Agentic architecture to create a supervised fine-tuning dataset; to train small but specialized language models specifically for crafting a threat response in the context of O-RAN 5G.
@@
    Currently, we supply network facts like CU configurations and parameters at inference time, which makes it slow and costly. Through supervised fine tuning, this information can be baked into the model parameters. We also plan to expand the use case to more than 5 attacks that we currently support, with tools designed to address a wide range of threats.
@@ IV-B Accessible and robust O-RAN security
We plan to integrate MobiLLM into real-world 6G deployments. As initiatives like the U.S. government’s Open Centralized Unit Distributed Unit (OCUDU) project accelerate the creation of open 6G platforms, MobiLLM can be embedded as a core security layer to not only enhance network resilience and trust from the outset, but also cut down network operator training cost. We plan to develop more advanced security guardrails to protect MobiLLM from adversarial attacks like backdoor attacks [16] and prompt injection attacks [14]. This involves creating stricter, verifiable rules for its API interactions to defend against threats like prompt injection and ensure its operational integrity in mission-critical environments.
@@ Acknowledgements
This research was supported by a Small Business Innovation Research (SBIR) Phase I award N6893625C0023 from the Naval Air Warfare Center (NAWC), the NSF convergence accelerator program under award ITE-2326882, and the Google Cloud Research Credits Program.